. For the first five hours of a flight, a plane is assisted by a 100 mile per hour wind. For the next three hours, it is hindered by a 100 mile per hour wind. If the total distance traveled is 3400 miles, how fast is the plane in windless conditions? Please show me the steps.
A. 500 mph B. 400 mph C. 300 mph D. 200 mph
E.None

Answer by lwsshak3(11628)

For the first five hours of a flight, a plane is assisted by a 100 mile per hour wind. For the next three hours, it is hindered by a 100 mile per hour wind. If the total distance traveled is 3400 miles, how fast is the plane in windless conditions? Please show me the steps.
A. 500 mph B. 400 mph C. 300 mph D. 200 mph
E.None
**
let x=speed of plane in windless conditions
(x+100)=speed of plane with wind
(x-100)=speed of plane against wind
distance=speed*travel time
..
5(x+100)+3(x-100)=3400
5x+500+3x-300=3400
8x=3200
x=400 mph
ans: B